# Primes and The Field of Values of Characters

### v1

@@ 1. Introduction
Next we turn to the related problem of classifying the fields of values of the irreducible characters of finite groups of order not divisible by $p$. This leads us to the following basic problem: is it always true that ${\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $|G|$? Interestingly, the recently proposed inductive Feit condition ([BKNT25]) implies what seemed to be an intractable problem.
@@
    If $G$ is a finite group and $\chi\in{\rm Irr}(G)$, then ${\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $|G|$.
@@
    Let $\mathbb{F}/{\mathbb{Q}}$ be an abelian extension and let $p$ be a prime. Then $\mathbb{F}={\mathbb{Q}}(\chi)$ for some $\chi\in{\rm Irr}(G)$, where $G$ is a finite group and $p$ does not divide $|G|$, if and only if both $c(\mathbb{F})$ and ${\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}]$ are not divisible by $p$.
@@ 3. Theorem D
Since the Harish-Chandra induction $R^{G}_{L_{\chi}}$ commutes with Galois action, it suffices to prove the claim in the case $m=1$. Now, if $d_{1}>1$, then $\chi=S(s_{1},\lambda_{1})$ is Lusztig induced from the character $S(s_{1},\lambda_{1})$ of the Levi subgroup $L=\operatorname{GL}_{k_{1}}(q^{d_{1}})$ of $G=\operatorname{GL}_{k_{1}d_{1}}(q)$, see [FS82, p. 116]. Since the Lefschetz numbers are all integers (see e.g. [DM91, p. 140]), Green functions involved in the Lusztig induction $R^{G}_{L}$ rational-valued, and thus $R^{G}_{L}$ also commutes with Galois actions. Thus it suffices to prove the claim in the case $m=d_{1}=1$. In the latter case, $\chi=S(1,\lambda_{1})\hat{s_{1}}$, where $S(1,\lambda_{1})$ is a rational (unipotent) character, and $\hat{s_{1}}$ is a linear character of $G/[G,G]$ of order equal to the order of $s_{1}$, see [FS82, (1.16)], so we are done. ∎
@@
Theorem D follows from Theorems 3.2, 3.3, and 3.6.
@@ 4. Theorem E
    If $G$ is a finite group and $\chi\in{\rm Irr}(G)$, then ${\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $|G|$.
@@
    Let $\mathbb{F}/{\mathbb{Q}}$ be an abelian extension and let $p$ be a prime. Then $\mathbb{F}={\mathbb{Q}}(\chi)$ for some $\chi\in{\rm Irr}(G)$, where $G$ is a finite group and $p$ does not divide $|G|$, if and only if both $c(\mathbb{F})$ and ${\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}]$ are not divisible by $p$.
@@
We argue by induction on $|G|$. If $\chi$ is linear, then ${\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]=1$, and we are done. We may assume that $\chi$ is not linear. By Theorem 4.1, there exists $H<G$ and $\psi\in\operatorname{Irr}(H)$ such that ${\mathbb{Q}}(\chi)\subseteq{\mathbb{Q}}(\psi)$ and ${\mathbb{Q}}(\psi):{\mathbb{Q}}(\chi)]$ divides $|G:H|$. By induction, we have that ${\mathbb{Q}}_{c(\psi)}:{\mathbb{Q}}(\psi)]$ divides $|H|$. Since ${\mathbb{Q}}(\chi)\subseteq{\mathbb{Q}}(\psi)$, we have that $c(\chi)$ divides $c(\psi)$. Let $K={\mathbb{Q}}_{c(\chi)}\cap{\mathbb{Q}}(\psi)$. By elementary Galois theory, ${\mathbb{Q}}_{c(\chi)}:K]$ divides ${\mathbb{Q}}_{c(\psi)}:{\mathbb{Q}}(\psi)]$, which divides $|H|$. Also, $K:{\mathbb{Q}}(\chi)]$ divides ${\mathbb{Q}}(\psi):{\mathbb{Q}}(\chi)]$, which divides $|G:H|$, and part (a) follows.
Now, let $\mathbb{F}/{\mathbb{Q}}$ be an abelian extension and let $p$ be a prime. By Lemma 2.3, there exists a group $G$ of order $c(\mathbb{F})[{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}]$ and $\chi\in\operatorname{Irr}(G)$ such that ${\mathbb{Q}}(\chi)=F$. Hence, if $c(\mathbb{F})[{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}]$ is not divisible by $p$, we have that $G$ is a $p^{\prime}$-group. Conversely, assume that $G$ is a $p^{\prime}$-group and $\chi\in{\rm Irr}(G)$. Since $c(\chi)$ divides $|G|$, we have that $c(\chi)$ is not divisible by $p$. By part (a), we have that ${\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $|G|$, and we are done. ∎

### v2

@@ 1. Introduction
  4. (d)
    $G$ is the special linear group $\operatorname{SL}_{n}(q)$ and $p\nmid\gcd(n,q-1)$.
  5. (e)
    $G$ is the special unitary group $\operatorname{SU}_{n}(q)$ and $p\nmid\gcd(n,q+1)$.
Also see Theorem 3.9 for results concerning $p^{\prime}$-extensions of $\operatorname{SL}_{n}(q)$ and $\operatorname{SU}_{n}(q)$.
Next we turn to the related problem of classifying the fields of values of the irreducible characters of finite groups whose order has a given $p$-part (in particular, groups of $p^{\prime}$-order). This leads us to the following basic problem: is it always true that $c(\chi)[{\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $|G|$? Interestingly, the recently proposed inductive Feit condition ([BKNT25]) implies what seemed to be an intractable problem.
@@
    If $G$ is a finite group and $\chi\in{\rm Irr}(G)$, then $c(\chi)[{\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $|G|$.
@@
    Let $\mathbb{F}/{\mathbb{Q}}$ be an abelian extension, let $p$ be a prime, and $b\in{\mathbb{Z}}_{\geq 0}$. Then $\mathbb{F}={\mathbb{Q}}(\chi)$ for some $\chi\in{\rm Irr}(G)$, where $G$ is a finite group with $\nu_{p}(|G|)=b$, if and only if
    $\nu_{p}(c(\mathbb{F})[{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}])\leq b.$
@@ 3. Theorem D
Since the Harish-Chandra induction $R^{G}_{L_{\chi}}$ commutes with Galois action, it suffices to prove the claim in the case $m=1$. Now, if $d_{1}>1$, then $\chi=S(s_{1},\lambda_{1})$ is Lusztig induced from the character $S(s_{1},\lambda_{1})$ of the Levi subgroup $L=\operatorname{GL}_{k_{1}}(q^{d_{1}})$ of $G=\operatorname{GL}_{k_{1}d_{1}}(q)$, see [FS82, p. 116]. Since the Lefschetz numbers are all integers (see e.g. [DM91, p. 140]), Green functions involved in the Lusztig induction $R^{G}_{L}$ are rational-valued, and thus $R^{G}_{L}$ also commutes with Galois actions. Thus it suffices to prove the claim in the case $m=d_{1}=1$. In the latter case, $\chi=S(1,\lambda_{1})\hat{s_{1}}$, where $S(1,\lambda_{1})$ is a rational (unipotent) character, and $\hat{s_{1}}$ is a linear character of $G/[G,G]$ of order equal to the order of $s_{1}$, see [FS82, (1.16)], so we are done. ∎
@@
We will need the following refinement of [NT21, Lemma 4.2]:
Proposition 3.7.
Let $p$ be a prime and let $N$ be a normal subgroup of a finite group $G$ with $p\nmid|G/N|$. Let $\theta\in\operatorname{Irr}(N)$, $J:=\mathrm{Stab}_{G}(\theta)$, and let $\mathbb{F}$ by any extension of $\mathbb{Q}$} inside ${\mathbb{Q}}_{|G|}$. Then the following statements hold.
  1. (i)
    Suppose that $\chi\in\operatorname{Irr}(G|\theta)$, $a=\nu_{p}(c(\chi))$, and $b=\nu_{p}(c(\theta))$. If $a,b\geq 1$ or $p=2$, then $a=b$. If $a=0$ then $b\leq 1$, and if $b=0$ then $a\leq 1$. Furthermore, $p\nmid[\mathbb{F}(\theta):(\mathbb{F}(\chi)\cap\mathbb{F}(\theta))]$. In fact, every prime divisor of $\mathbb{F}(\theta):(\mathbb{F}(\chi_{N})\cap\mathbb{F}(\theta))]$ divides $|G/N|$.
  2. (ii)
    Suppose that all the characters in $\operatorname{Irr}(J|\theta)$ have the same degree (this holds if, for instance, $J/N$ is cyclic). Then, there exists some $\chi\in\operatorname{Irr}(G|\theta)$ such that $p\nmid[\mathbb{F}(\chi):(\mathbb{F}(\chi)\cap\mathbb{F}(\theta))]$.
Proof.
(i) The first statement is just [NT21, Lemma 4.2(ii)]. The second statement follows from [H24, Lemma 2.4(ii)], where it was shown that if $\max\{a,b\}\geq 2$, then $a=b$.
For the third statement, by Clifford’s theorem we have $\chi_{N}=e\sum^{t}_{i=1}\theta_{i}$, where $t=[G:J]$ is coprime to $p$, and $\theta=\theta_{1},\theta_{2},\ldots,\theta_{t}$ are the $t$ distinct $G$-conjugates of $\theta$; in particular, ${\mathbb{Q}}(\theta_{i})={\mathbb{Q}}(\theta)$ for all $i$. Now let $\Gamma:=\mathrm{Gal}(L/K)$, where $L:=\mathbb{F}(\chi,\theta)$ and $K:=\mathbb{F}(\chi)$, and let $\Gamma_{p}$ denote the Sylow $p$-subgroup of $\Gamma$. Then, any $\sigma\in\Gamma_{p}$ fixes $\chi_{N}$, and so it permutes $\theta_{1},...,\theta_{t}$. Thus the $p$-group $\Gamma_{p}$ acts on the set $\{\theta_{1},\ldots,\theta_{t}\}$ of $p^{\prime}$-size, and hence $\Gamma_{p}$ fixes some $\theta_{j}$. Setting $M:=\mathbb{F}(\theta_{j})=\mathbb{F}(\theta)$, we see that $MK$ is contained in the fixed field $L^{\Gamma_{p}}$ which has $p^{\prime}$-degree $\Gamma:\Gamma_{p}]$ over $K$ by Galois correspondence. It follows that $p\nmid[MK:K]=[M:K\cap M]=[\mathbb{F}(\theta):(\mathbb{F}(\chi)\cap\mathbb{F}(\theta))]$.
The same proof using Sylow subgroups of $\mathrm{Gal}(\mathbb{F}(\theta,\chi_{N})/\mathbb{F}(\chi_{N}))$ yields the fourth statement.
(ii) The assumption implies that all the characters $\gamma$ in $\operatorname{Irr}(J|\theta)$ have the same ramification index $e$: $\gamma_{N}=e\theta$, and so, by the Clifford correspondence [I76, (6.11)], if $\operatorname{Irr}(G|\theta)=\{\chi_{1},\ldots,\chi_{s}\}$ then
$\theta^{G},\chi_{i}]_{G}=[\theta,(\chi_{i})_{N}]_{N}=e$
for all $i$. Hence,
(3.2) $\theta^{G}=e\sum^{s}_{i=1}\chi_{i},$
and thus
$\bigl(\theta^{G}\bigr)_{N}=e\sum^{s}_{i=1}(\chi_{i})_{N}=e^{2}s\sum^{t}_{j=1}\theta_{j}$
in the notation of (i). It follows that $e^{2}st=|G/N|$, and so $p\nmid s$.
Now let $\Gamma:=\mathrm{Gal}(L/K)$, where $L:=\mathbb{F}(\chi_{1},\ldots,\chi_{s},\theta)$ and $K:=\mathbb{F}(\theta)$, and let $\Gamma_{p}$ denote the Sylow $p$-subgroup of $\Gamma$. Then, any $\sigma\in\Gamma_{p}$ fixes $\theta^{G}$, and so it permutes $\chi_{1},...,\chi_{s}$ by (3.2). Thus the $p$-group $\Gamma_{p}$ acts on the set $\{\chi_{1},\ldots,\chi_{s}\}$ of $p^{\prime}$-size, and hence $\Gamma_{p}$ fixes some $\chi=\chi_{i}$. Setting $M:=\mathbb{F}(\chi)$, we see that $MK$ is contained in the fixed field $L^{\Gamma_{p}}$ which has $p^{\prime}$-degree $\Gamma:\Gamma_{p}]$ over $K$ by Galois correspondence. It follows that $p\nmid[MK:K]=[M:K\cap M]=[\mathbb{F}(\chi):(\mathbb{F}(\chi)\cap\mathbb{F}(\theta))]$. ∎
Now we can prove the following “going-up and going-down” result:
Theorem 3.8.
Let $p$ be a fixed prime, $G$ a finite group, and $N$ a normal subgroup of $G$ such that $p\nmid|G/N|$. Then the following statements hold.
  1. (i)
    Suppose Conjecture B holds for all irreducible characters of NN at the prime $p$. Then Conjecture B holds for all irreducible characters of $G$ at the prime $p$.
  2. (ii)
    Suppose that $G/N$ is solvable and that Conjecture B holds for all irreducible characters of $G$ at the prime $p$. Then Conjecture B holds for all irreducible characters of NN at the prime $p$.
Proof.
(i) It suffices to prove the theorem in the case $\chi\in\operatorname{Irr}(G)$ has $a:=\nu_{p}(c(\chi))\geq 2$. By Proposition 3.7(i), for any $\theta\in\operatorname{Irr}(N)$ lying under $\chi$ we have
$\nu_{p}(c(\theta))=a,$
furthermore,
$p\nmid[M:K]$
with $M:={\mathbb{Q}}_{p}(\theta)$ and $K:={\mathbb{Q}}_{p}(\chi)\cap{\mathbb{Q}}_{p}(\theta)$. Since
$=[{\mathbb{Q}}_{p^{a}}M:{\mathbb{Q}}_{p^{a}}]/[{\mathbb{Q}}_{p^{a}}K:{\mathbb{Q}}_{p^{a}}]$
$=[M:({\mathbb{Q}}_{p^{a}}\cap M)]/[K:({\mathbb{Q}}_{p^{a}}\cap K)]$
$=[M:K]/[({\mathbb{Q}}_{p^{a}}\cap M):({\mathbb{Q}}_{p^{a}}\cap K)]$
is an integer, it follows that
$p\nmid[({\mathbb{Q}}_{p^{a}}\cap{\mathbb{Q}}_{p}(\theta)):({\mathbb{Q}}_{p^{a}}\cap{\mathbb{Q}}_{p}(\chi)\cap{\mathbb{Q}}_{p}(\theta))].$
But ${\mathbb{Q}}_{p^{a}}\cap{\mathbb{Q}}_{p}(\chi)\cap{\mathbb{Q}}_{p}(\theta)\supseteq{\mathbb{Q}}_{p}$ and ${\mathbb{Q}}_{p^{a}}:{\mathbb{Q}}_{p}]=p^{a-1}$, so in fact we have shown that
${\mathbb{Q}}_{p^{a}}\cap{\mathbb{Q}}_{p}(\theta)\subseteq{\mathbb{Q}}_{p^{a}}\cap{\mathbb{Q}}_{p}(\chi).$
Denoting $p^{c}:=[{\mathbb{Q}}_{p^{a}}:({\mathbb{Q}}_{p^{a}}\cap{\mathbb{Q}}_{p}(\theta))]$ and $p^{d}:=[{\mathbb{Q}}_{p^{a}}:({\mathbb{Q}}_{p^{a}}\cap{\mathbb{Q}}_{p}(\chi))]$, we now have
(3.3) $c\geq d.$
By hypothesis, $p^{c}|\theta(1)$. Next, $p\nmid|G/N|$ implies that $p\nmid\chi(1)/\chi(1)$, and so $p^{c}|\chi(1)$. Hence $p^{d}|\chi(1)$ by (3.3), and we are done.
(ii) It suffices to prove the theorem in the case $G/N$ is cyclic, and furthermore $\theta\in\operatorname{Irr}(N)$ has $b=\nu_{p}(c(\theta))\geq 2$. By Proposition 3.7(i), for any $\chi\in\operatorname{Irr}(G|\theta)$ we have
$\nu_{p}(c(\chi))=b.$
Since $\mathrm{Stab}_{G}(\theta)/N$ is cyclic in our case, using Proposition 3.7(ii) with $\mathbb{F}:={\mathbb{Q}}_{p}$, there exists some $\chi\in\operatorname{Irr}(G|\theta)$ such that
$p\nmid[{\mathbb{Q}}_{p}(\chi):({\mathbb{Q}}_{p}(\chi)\cap{\mathbb{Q}}_{p}(\theta))].$
Arguing as in (i), we see that
$p\nmid[({\mathbb{Q}}_{p^{b}}\cap{\mathbb{Q}}_{p}(\chi)):({\mathbb{Q}}_{p^{b}}\cap{\mathbb{Q}}_{p}(\chi)\cap{\mathbb{Q}}_{p}(\theta))],$
and hence
${\mathbb{Q}}_{p^{b}}\cap{\mathbb{Q}}_{p}(\chi)\subseteq{\mathbb{Q}}_{p^{b}}\cap{\mathbb{Q}}_{p}(\theta).$
Denoting $p^{c}:=[{\mathbb{Q}}_{p^{b}}:({\mathbb{Q}}_{p^{b}}\cap{\mathbb{Q}}_{p}(\chi))]$ and $p^{d}:=[{\mathbb{Q}}_{p^{b}}:({\mathbb{Q}}_{p^{b}}\cap{\mathbb{Q}}_{p}(\theta))]$, we now have
(3.4) $c\geq d.$
By hypothesis, $p^{c}|\chi(1)$. Next, $p\nmid|G/N|$ implies that $p\nmid\chi(1)/\theta(1)$, and so $p^{c}|\theta(1)$. Hence $p^{d}|\theta(1)$ by (3.4), and we are done. ∎
Theorem 3.9.
The following statements hold.
  1. (i)
    Suppose that $S=\operatorname{SL}_{n}(q)$ and $p\nmid\gcd(n,q-1)$. Then Theorem D at the prime $p$ holds for $S=\operatorname{SL}_{n}(q)$ and for any finite group $L\rhd S$ with $p\nmid|L/S|$.
  2. (ii)
    Suppose that $S=\operatorname{SU}_{n}(q)$ and $p\nmid\gcd(n,q+1)$. Then Theorem D at the prime $p$ holds for $S=\operatorname{SU}_{n}(q)$ and for any finite group $L\rhd S$ with $p\nmid|L/S|$.
Proof.
(i) Embed $S$ in $G:=\operatorname{GL}_{n}(q)$, and consider $H:={\mathbf{Z}}(G)S$. Since $G/H\cong\mathsf{C}_{\gcd(n,q-1)}$, by Theorems 3.6 and 3.8(ii), Conjecture B holds for $H$ at $p$. Note that
$H={\mathbf{O}}_{p^{\prime}}({\mathbf{Z}}(G))S\times{\mathbf{O}}_{p}({\mathbf{Z}}(G))$
is a direct product. Indeed, suppose that $\alpha\cdot\mathrm{Id}=\beta g$ where $\alpha\cdot\mathrm{Id}\in{\mathbf{O}}_{p}({\mathbf{Z}}(G))$, $\beta\cdot\mathrm{Id}\in{\mathbf{O}}_{p^{\prime}}({\mathbf{Z}}(G))$, and $g\in S$. So $\alpha\in\mathbb{F}_{q}^{\times}$ is a $p$-element and $\beta\in\mathbb{F}_{q}^{\times}$ is a $p^{\prime}$-element. On the other hand,
$\alpha^{n}={\rm det}(\alpha\cdot\mathrm{Id})={\rm det}(\beta g)=\beta^{n},$
so $\alpha^{n}$ is both a $p$-element and a $p^{\prime}$-element, whence $\alpha^{n}=1$. But $\alpha^{q-1}=1$, so the order of the $p$-element $\alpha$ divides $\gcd(n,q-1)$. As $p\nmid\gcd(n,q-1)$, we conclude $\alpha=1$.
As ${\mathbf{O}}_{p^{\prime}}({\mathbf{Z}}(G))S$ is a direct factor of $H$, Conjecture B holds for ${\mathbf{O}}_{p^{\prime}}({\mathbf{Z}}(G))S$. Since $S$ is a normal subgroup of ${\mathbf{O}}_{p^{\prime}}({\mathbf{Z}}(G))S$ with quotient a cyclic $p^{\prime}$-group, so Conjecture B holds for $S$ at $p$ by Theorem 3.8(ii). By Theorem 3.8(i), Conjecture B then holds for $L$ at $p$.
(ii) The arguments are entirely similar to the ones in (i). ∎
Proof of Theorem D.
It follows from Theorems 3.2, 3.3, 3.6, and 3.9. ∎
@@ 4. Theorem E
    If $G$ is a finite group and $\chi\in{\rm Irr}(G)$, then $c(\chi)[{\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $|G|$.
@@
    Let $\mathbb{F}/{\mathbb{Q}}$ be an abelian extension, let $p$ be a prime, and $b\in{\mathbb{Z}}_{\geq 0}$. Then $\mathbb{F}={\mathbb{Q}}(\chi)$ for some $\chi\in{\rm Irr}(G)$, where $G$ is a finite group with $\nu_{p}(|G|)=b$, if and only if
    $\nu_{p}(c(\mathbb{F})[{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}])\leq b.$
@@
We argue by induction on $|G|$. If $\chi$ is linear, then $c(\chi)[{\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]=c(\chi)$ divides the exponent of $G$, and we are done. We may assume that $\chi$ is not linear. By Theorem 4.1, there exists $H<G$ and $\psi\in\operatorname{Irr}(H)$ such that ${\mathbb{Q}}(\chi)\subseteq{\mathbb{Q}}(\psi)$ and ${\mathbb{Q}}(\psi):{\mathbb{Q}}(\chi)]$ divides $|G:H|$. By induction, we have that $c(\psi)[{\mathbb{Q}}_{c(\psi)}:{\mathbb{Q}}(\psi)]$ divides $|H|$. Since ${\mathbb{Q}}(\chi)\subseteq{\mathbb{Q}}(\psi)$, we have that $c(\chi)$ divides $c(\psi)$. Let $K={\mathbb{Q}}_{c(\chi)}\cap{\mathbb{Q}}(\psi)$. By elementary Galois theory, $c(\chi)[{\mathbb{Q}}_{c(\chi)}:K]$ divides $c(\psi)[{\mathbb{Q}}_{c(\psi)}:{\mathbb{Q}}(\psi)]$, which divides $|H|$. Also, $K:{\mathbb{Q}}(\chi)]$ divides ${\mathbb{Q}}(\psi):{\mathbb{Q}}(\chi)]$, which divides $|G:H|$, and part (a) follows.
Now, let $\mathbb{F}/{\mathbb{Q}}$ be an abelian extension, let $p$ be a prime, and $b\in{\mathbb{Z}}_{\geq 0}$ such that $\nu_{p}(c(\mathbb{F})[{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}])\leq b$. By Lemma 2.3, there exists a group $M$ of order $c(\mathbb{F})[{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}]$ and $\varphi\in\operatorname{Irr}(M)$ such that ${\mathbb{Q}}(\varphi)=F$. The direct product $G:=M\times C$, where $C$ is the cyclic $p$-group of order $p^{b}/(c(\mathbb{F})_{p}[{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}]]_{p})$, and the character $\chi:=\varphi\otimes\textbf{1}_{C}\in\operatorname{Irr}(G)$ satisfy the required condition. Conversely, assume that $G$ is a finite group such that $|G|_{p}=p^{b}$ and $\chi\in{\rm Irr}(G)$. By part (a), we have that $c(\chi)[{\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $|G|$, which implies that $c(\chi)_{p}[{\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]_{p}\leq|G|_{p}=p^{b}$, and we are done. ∎